OpenStudy (anonymous):

Solve for x: -3(x + 3) = -3(x + 1) - 5 Answer -6 No solution All real numbers 6

jimthompson5910 (jim_thompson5910):

-3(x+3)=-3(x+1)-5 -3x-9=-3x-3-5 -3x-9=-3x-8 -3x=-3x-8+9 -3x+3x=-8+9 0x=-8+9 0x=1 0=1 Since this equation is NEVER true for any x value, this means that there are no solutions.